# Scattering of the defocusing Calogero–Moser derivative nonlinear Schrödinger equation

### v2

@@ Abstract
Using the Gérard-type explicit formula, we prove the scattering result of solutions to this equation with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. We also characterize the scattering term using the distorted Fourier transform associated with the Lax operator. Following our approach developed in this paper, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases.
@@
Acknowledgment The author is currently a postdoctoral researcher at the Department of Mathematics and Computer Science, University of Basel. Part of this work was carried out during his PhD studies at the Laboratoire de Mathématiques d’Orsay (UMR 8628), Université Paris-Saclay. The author is deeply grateful to Patrick Gérard for his suggestions throughout this work.
@@ 1 Introduction
They proved that if $u\in H_{+}^{s}(\mathbb{R}):=H^{s}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ with $s$ large enough, then equation (1.3) admits a Lax pair structure and has infinitely many conservation laws. Using these conservation laws, they established global well-posedness of (1.3) in $H_{+}^{s}(\mathbb{R})$ for all $s>\tfrac{1}{2}$ when the initial data has small mass ($\|u_{0}\|_{L^{2}}^{2}<2\pi$). Later, Killip, Laurens, and Vişan combined the explicit formulae of (1.1) and (1.3) with methods based on commuting flows to show global well-posedness of both equations in $H_{+}^{s}(\mathbb{R})$ for $s\geq 0$, under the assumption of small mass initial data in the focusing case [17]. Hogan and Kowalski showed that for every $\varepsilon>0$, one can find initial data $u_{0}\in H_{+}^{\infty}$ with mass $2\pi+\varepsilon$ such that the corresponding maximal lifespan solution $u:(T_{-},T_{+})\times\mathbb{R}\to\mathbb{C}$ satisfies $\lim_{t\to T_{\pm}}\|u(t)\|_{H^{s}}=\infty$ for all $s>0$ [16]. K. Kim, T. Kim and Kwon constructed finite-time blow-up solutions in $\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ to (1.3) with mass arbitrarily close to $2\pi$ [18]. More recently, the author showed global well-posedness of (1.1) with a non-vanishing condition at infinity and also established Gérard-type explicit formulae for the corresponding solutions [9].
In [10], Frank and Read proposed an inverse scattering transform for the focusing (CM-DNLS) (1.3) using the distorted Fourier transform. Furthermore, Badreddine characterized the zero dispersion limit of solutions to (1.1) and (1.3) with initial data $u_{0}\in L_{+}^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ through Gérard-type explicit formulae, and identified this limit in terms of the branches of the multivalued solution to the inviscid Burgers–Hopf equation [5]. She also established global well-posedness of (1.1) and (1.3) on the circle in $H_{+}^{s}(\mathbb{T})$ for any $s\geq 0$, under the assumption of small mass initial data in the focusing case [3], and further classified the traveling wave solutions of (1.1) and (1.3) on the circle [4]. Notice that a new numerical scheme based on explicit formulae was recently developped for (1.1) and (1.3) on the circle [2].
@@
In this paper, we applied the Gérard-type explicit formula to prove the scattering result for solutions to the defocusing (CM-DNLS) (1.1) with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. Our analysis also relies on the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$, and we can characterize the scattering term using this distorted Fourier transform. Following our approach developed in this paper, combined with the distorted Fourier transform corresponding to the focusing (CM-DNLS) (1.3) introduced in [10, Section 4], we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) (1.3) in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$.
@@ 1.1 The Lax pair
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\left|\widehat{f}\left(0_{+}\right)\right|^{2}=-4\pi\operatorname{Im}\left\langle X^{*}f\mid f\right\rangle\leq 4\pi\|X^{*}f\|_{L^{2}}\|f\|_{L^{2}}.$
@@ 1.3 Gérard-type explicit formula
([17, Theorem 1.7]) Let $u\in C\left(\mathbb{R},L_{+}^{2}(\mathbb{R})\right)$ be the solution to (1.1) with $u(0)=u_{0}$. Then for any $t\in\mathbb{R}$,
$\forall z\in\mathbb{C}_{+},\quad u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right].$ (1.8)
@@
Now we introduce the following lemma concerning the Gérard-type explicit formula, which will be useful in the proof of our main theorem later on.
Lemma 1.4.
Given $u_{0}\in H_{+}^{1}(\mathbb{R})$ and $f_{0}\in L_{+}^{2}(\mathbb{R})$, and we define
$f(t,z):=\frac{1}{2i\pi}I_{+}\left(\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}f_{0}\right),\qquad\forall z\in\mathbb{C}_{+},\,\forall t\in\mathbb{R}.$
Then we have
$|f(t,z)|\leq C(z)\|f\|_{L^{2}(\mathbb{R})}\leq C(z)\|f_{0}\|_{L^{2}(\mathbb{R})},\quad\forall z\in\mathbb{C}_{+},\,\forall t\in\mathbb{R},$ (1.9)
where $C(z)$ depends only on $z$.
Proof.
In [8, Lemma 3.1], the author proved an analogous result corresponding to the Benjamin–Ono equation on the line. One can follow the approach in the proof of [8, Lemma 3.1] to show Lemma 1.4. ∎
@@ 1.4 Main result
In this paper, we used the Gérard-type explicit formula (1.8) to study the long-time behavior of (1.1). More precisely, we proved the following scattering result for solutions to (1.1) with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u_{0}\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u_{0}\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$.
@@
Let $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{w\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. Then the solution $u\in C(\mathbb{R};H_{+}^{1}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
@@
As mentioned in the introduction, Kim and Kwon proved the scattering result for solutions to (1.3) in $H^{1,1}(\mathbb{R})$ [19]. More precisely, their appraoch relies on the pseudo-conformal symmetry, and thus requires $xu_{0}\in L^{2}(\mathbb{R})$. Also, it is expected that we can follow their approach to show the scattering result for the defocusing (CM-DNLS) (1.1) in $H^{1,1}(\mathbb{R})$. The great progress we made in this paper is to extend the chiral scattering result to more singular chiral initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. On the other hand, we can characterize the scattering term using the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$ of the defocusing (CM-DNLS) (1.1). In fact, following the method used in this paper, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [10, Section 4], we can also establish the scattering result for solutions to the focusing (CM-DNLS) (1.3) with initial data in $H_{+}^{1,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ for $\alpha>\frac{1}{4}$, which is an extension of the chiral scattering result for the focusing (CM-DNLS) (1.3) in [19]. See Remark 3.2 for details. Another highlight of this paper is that it is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. See also [6] for the study on the long-time behavior of the Benjamin–Ono equation with rational initial data and [12] for the study on the long-time behavior of the half-wave maps equation with rational initial data.
@@
In [19], Kim and Kwon also showed the soliton resolution for global solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R})$. With a slight modification of the proof of Theorem 1.5, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [10, Section 4], we can conclude the asymptotic bound-state/radiation decomposition for global solutions to (1.3) in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$, i.e., if $u\in C(\mathbb{R};H_{+}^{1}(\mathbb{R}))$ is a global solution to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}\in H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$, then
@@
in $L_{+}^{2}(\mathbb{R})$. Here $v_{0}$ can be defined as in Theorem 1.5, using the distorted Fourier transform introduced in [10, Section 4]. To establish soliton resolution in this setting, one would need to show that each bound state asymptotically converges to a traveling soliton profile. See Remark 3.2 and Remark 3.3 for details.
@@
One can observe that the (unique) ground state solution $\frac{\sqrt{2}}{x+i}$ to the focusing (CM-DNLS) (1.3) (see [11, Lemma 4.1]) is in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\frac{1}{4}<\alpha<\frac{1}{2}$ but not in $H^{1,1}(\mathbb{R})$. This shows that our progress on the long-time analysis of (1.1) and (1.3) is crucial.
@@ 2 Preliminaries
In this section, we study the generalized eigenfunctions and the distorted Fourier transform associated with the Lax operator $L_{u}$ of (1.1). For $u\in H_{+}^{1}(\mathbb{R})$, it is known that $L_{u}$ has no eigenvalue in this case. Under some additional decay assumption on $u\in H_{+}^{1}(\mathbb{R})$, we show the existence of the generalized eigenfunctions in $L_{+}^{\infty}(\mathbb{R})$ and then introduce the corresponding distorted Fourier transform. See also [10, Section 4] for the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3).

### v3

@@ Abstract
Using the Gérard-type explicit formula, we prove the scattering result of solutions to this equation with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. We also characterize the scattering term using the distorted Fourier transform associated with the Lax operator. Following our approach developed in this paper, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases.
@@
Acknowledgment The author is currently a postdoctoral researcher at the Department of Mathematics and Computer Science, University of Basel. Part of this work was carried out during his PhD studies at the Laboratoire de Mathématiques d’Orsay (UMR 8628), Université Paris-Saclay. The author is deeply grateful to Patrick Gérard for his insightful comments throughout this work. The author also wishes to thank Thierry Laurens for his helpful suggestions.
@@ 1 Introduction
They proved that if $u\in H_{+}^{s}(\mathbb{R}):=H^{s}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ with $s$ large enough, then the equation (1.3) admits a Lax pair structure and has infinitely many conservation laws. Using these conservation laws, they established global well-posedness of (1.3) in $H_{+}^{s}(\mathbb{R})$ for all $s>\tfrac{1}{2}$ when the initial data has small mass ($\|u_{0}\|_{L^{2}}^{2}<2\pi$). Later, Killip, Laurens, and Vişan combined the explicit formulae of (1.1) and (1.3) with methods based on commuting flows to show global well-posedness of both equations in $H_{+}^{s}(\mathbb{R})$ for $s\geq 0$, under the assumption of small mass initial data in the focusing case [19]. Hogan and Kowalski showed that for every $\varepsilon>0$, one can find initial data $u_{0}\in H_{+}^{\infty}$ with mass $2\pi+\varepsilon$ such that the corresponding maximal lifespan solution $u:(T_{-},T_{+})\times\mathbb{R}\to\mathbb{C}$ satisfies $\lim_{t\to T_{\pm}}\|u(t)\|_{H^{s}}=\infty$ for all $s>0$ [17]. K. Kim, T. Kim and Kwon constructed finite-time blow-up solutions in $\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ to (1.3) with mass arbitrarily close to $2\pi$ [20]. The quantized blow-up dynamics for the focusing (CM-DNLS) (1.3) was studied in [18]. More recently, the author showed global well-posedness of (1.1) with a non-vanishing condition at infinity and also established Gérard-type explicit formulae for the corresponding solutions [10].
In [22], Matsuno studied the Cauchy problem for the defocusing (CM-DNLS) (1.1) with a non-zero constant condition at infinity, developing an inverse scattering transform formulation. In [11], Frank and Read proposed an inverse scattering transform for the focusing (CM-DNLS) (1.3) using the distorted Fourier transform.
Furthermore, Badreddine characterized the zero dispersion limit of solutions to (1.1) and (1.3) with initial data $u_{0}\in L_{+}^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ through Gérard-type explicit formulae, and identified this limit in terms of the branches of the multivalued solution to the inviscid Burgers–Hopf equation [5]. She also established global well-posedness of (1.1) and (1.3) on the circle in $H_{+}^{s}(\mathbb{T})$ for any $s\geq 0$, under the assumption of small mass initial data in the focusing case [3], and further classified the traveling wave solutions of (1.1) and (1.3) on the circle [4]. Notice that a new numerical scheme based on explicit formulae was recently developped for (1.1) and (1.3) on the circle [2]. More studies on the intermediate nonlinear Schrödinger equation can be found in [26, 27, 28, 8].
@@
In this paper, we applied the Gérard-type explicit formula to prove the scattering result for solutions to the defocusing (CM-DNLS) (1.1) with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. Our analysis also relies on the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$, and we can characterize the scattering term using this distorted Fourier transform. Following our approach developed in this paper, combined with the distorted Fourier transform corresponding to the focusing (CM-DNLS) (1.3) introduced in [11, Section 4], we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) (1.3) in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$.
@@ 1.1 The Lax pair
For $u\in L_{+}^{2}(\mathbb{R})$, the Lax operator $L_{u}$ can be defined via an approach using quadratic forms with $\operatorname{Dom}(L_{u})\subset H_{+}^{\frac{1}{2}}(\mathbb{R})$ (see also [12, Appendix A]).
For $u\in H_{+}^{2}(\mathbb{R})$, another operator in the Lax pair, denoted by $B_{u}$, is defined by
$B_{u}=-uT_{\partial_{x}\bar{u}}+\partial_{x}uT_{\bar{u}}+i\left(uT_{\bar{u}}\right)^{2}.$
@@
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\left|\widehat{f}\left(0_{+}\right)\right|^{2}=-4\pi\operatorname{Im}\left\langle X^{*}f,f\right\rangle_{L^{2}(\mathbb{R})}\leq 4\pi\|X^{*}f\|_{L^{2}}\|f\|_{L^{2}}.$
@@
1.2 Maximally dissipative operators and the Kato-Rellich theorem
In this paper, we will also apply the Miyadera–Gustafson theorem to show that operators remain maximally dissipative after some perturbations. To present the Miyadera–Gustafson theorem for maximally dissipative operators, we first need to introduce the following definition of the dissipative and maximally dissipative operators in Banach spaces.
Definition 1.1.
Let $(D(A),A)$ be an operator in a Banach space $\mathscr{B}$.
1. We say that $A$ is dissipative if for all $g\in D(A)$ and all $\lambda>0$,
$\|(\lambda I-A)g\|\geq\lambda\|g\|.$
2. We say that $A$ is maximal dissipative if it is dissipative and for all $h\in\mathscr{B}$ and for all $\lambda>0$, there exists $g\in D(A)$ such that $(\lambda I-A)g=h$.
Remark 1.2.
In fact, an operator $(D(A),A)$ in a Hilbert space $\mathscr{H}$ is dissipative if and only if for all $g\in D(A),\,\Re\langle Ag,g\rangle\leq 0$.
Remark 1.3.
Let $(D(A),A)$ be a maximally dissipative operator in a Banach space $\mathscr{B}$. From the definition of maximally dissipative operators, we can deduce that, for all $\lambda>0$, we have
$\|(\lambda I-A)^{-1}\|_{\mathscr{L}(\mathscr{B})}\leq\frac{1}{\lambda},\qquad\|A(\lambda I-A)^{-1}\|_{\mathscr{L}(\mathscr{B})}\leq 2.$ (1.8)
Since the Miyadera–Gustafson theorem involves the related notion of the perturbation of operators, we give the following definition of the relative bound of an operator with respect to another operator.
Definition 1.4.
Let $\left(D(A),A\right)$ and $\left(D(B),B\right)$ be densely defined linear operators on a Banach space $\mathscr{B}$. Suppose that:
(i) $D(A)\subset D(B)$;
(ii) For some $a$ and $b$ in $\mathbb{R}$ and all $\varphi\in D(A)$,
$\|B\varphi\|\leq a\|A\varphi\|+b\|\varphi\|.$
Then $B$ is said to be $A$-bounded. The infimum of such $a$ is called the relative bound of $B$ with respect to $A$. If the relative bound is 0, we say that $B$ is infinitesimally small with respect to $A$.
Then we state the Miyadera–Gustafson theorem for maximally dissipative operators.
Theorem 1.5 (Miyadera–Gustafson theorem).
Let $\left(D(A),A\right)$ be a maximally dissipative operator which is densely defined on a Banach space $\mathscr{B}$ and assume $\left(D(B),B\right)$ to be dissipative and $A$-bounded with the relative bound smaller than 1. Then $\left(D(A),A+B\right)$ is also a maximally dissipative operator.
We refer to [25] for the proof of the Miyadera–Gustafson theorem.
@@ 1.4 Gérard-type explicit formula
([19, Theorem 1.7]) Let $u\in C\left(\mathbb{R},L_{+}^{2}(\mathbb{R})\right)$ be the solution to (1.1) with $u(0)=u_{0}$. Then for any $t\in\mathbb{R}$,
$\forall z\in\mathbb{C}_{+}:=\{z\in\mathbb{C}:\,\Im(z)>0\},\quad u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right].$ (1.9)
@@ 1.5 Main result
In this paper, we used the Gérard-type explicit formula (1.9) to study the long-time behavior of (1.1). More precisely, we proved the following scattering result for solutions to (1.1) with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u_{0}\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u_{0}\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$.
@@
Let $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R}):=\{w\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. Then the solution $u\in C(\mathbb{R};L_{+}^{2}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
@@
As mentioned in the introduction, Kim and Kwon proved the scattering result for solutions to (1.3) in $H^{1,1}(\mathbb{R})$ [21]. More precisely, their appraoch relies on the pseudo-conformal symmetry, and thus requires $xu_{0}\in L^{2}(\mathbb{R})$. Also, it is expected that we can follow their approach to show the scattering result for the defocusing (CM-DNLS) (1.1) in $H^{1,1}(\mathbb{R})$. The great progress we made in this paper is to extend the chiral scattering result to more singular chiral initial data in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$. On the other hand, we can characterize the scattering term using the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$ of the defocusing (CM-DNLS) (1.1). In fact, following the method used in this paper, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [11, Section 4], we can also establish the scattering result for solutions to the focusing (CM-DNLS) (1.3) with initial data in $L_{+}^{2,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ with some $\alpha>0$, which is an extension of the chiral scattering result for the focusing (CM-DNLS) (1.3) in [21]. See Remark 3.2 for details. Another highlight of this paper is that it is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. See also [6] for the study on the long-time behavior of the Benjamin–Ono equation with rational initial data and [13] for the study on the long-time behavior of the half-wave maps equation with rational initial data.
@@
In [21], Kim and Kwon also showed the soliton resolution for global solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R})$. With a slight modification of the proof of Theorem 1.9, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [11, Section 4], we can conclude the asymptotic bound-state/radiation decomposition for global solutions to (1.3) in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$, i.e., if $u\in C(\mathbb{R};L_{+}^{2}(\mathbb{R}))$ is a global solution to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ for some $\alpha>0$, then
@@
in $L_{+}^{2}(\mathbb{R})$. Here $v_{0}$ can be defined as in Theorem 1.9, using the distorted Fourier transform introduced in [11, Section 4]. To establish soliton resolution in this setting, one would need to show that each bound state asymptotically converges to a traveling soliton profile. See Remark 3.2 and Remark 3.3 for details.
@@
One can observe that the (unique) ground state solution $\frac{\sqrt{2}}{x+i}$ to the focusing (CM-DNLS) (1.3) (see [12, Lemma 4.1]) is in $L_{+}^{2,\alpha}(\mathbb{R})$ for $0<\alpha<\frac{1}{2}$ but not in $H^{1,1}(\mathbb{R})$. This shows that our progress on the long-time analysis of (1.1) and (1.3) is crucial.
@@ 2 Preliminaries
In this section, we study the generalized eigenfunctions and the distorted Fourier transform associated with the Lax operator $L_{u}$ of (1.1). For $u\in L_{+}^{2}(\mathbb{R})$, it is known that $L_{u}$ has no eigenvalue in this case. Under some additional decay assumption on $u\in L_{+}^{2}(\mathbb{R})$, we show the existence of the generalized eigenfunctions in $L_{+}^{\infty}(\mathbb{R})$ and then introduce the corresponding distorted Fourier transform. See also [11, Section 4] for the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3).